Calculate the total pressure in a 10 litre cylinder which contains 0.4 g He, 1.6g oxygen & 1.4 g of nitrogen at 300k assume ideal behavior for gases. A. 0.492 atmp B. 0.246 atmp C. 0.692 atmpD. Non of these

Grade:12

1 Answers

Arun
6 years ago
PV = nRT V = 10 litre T = 27ºC = 300K
nHe = 0.4/4 = 0.1
 N2 = 1.4/ 28= 0.05, 
O2 = 1.6/32= 0.05
Total no. of moles = 0.1 + 0.05 + 0.05  0.2
PV = nRT  P× 10 = 0.2 × 0.082 × 300 = 0.04926  P = 0.492 atm
Partial pressure = Total pressure × molefraction
PHe = 0.492 × 0.1/ 0.2
 = 0.246 atm

PN2 = 0.492 × 0.05/ 0.2
 = 0.123 atm

PO2 = 0.492 × 0.05/0.2
 = 0.123 atm